The sum of 3 consecutive odd integers is k. In terms of k, what is the sum of the 2 smaller of these integers?
Answer by Alan3354(69443)

k = n + n+2 + n+4
k = 3n + 6
n = k/3 - 2
-----------
n + n+2 = 2n + 2
= 2(k/3 - 2) + 2
= 2k/3 - 4 + 2
= 2k/3 - 2
or (2k - 6)/3